OpenStudy (anonymous):

How do you put 2(x-11)^2+4 in standard form?

OpenStudy (anonymous):

can anyone help?

OpenStudy (haseeb96):

just first open the bracket value by this formula then multiply by 2 and then add the constant to 4 (a-b)^2=a^2 -2ab + b^2

OpenStudy (xapproachesinfinity):

Expand to (x-11)^2 just like he said then distribute 2 continue....

OpenStudy (coconutjj):

2x^2-44x=-246